# Compatible $E$-differential forms on Lie algebroids over (pre-)multisymplectic manifolds

### v3

@@
Compatible $E$-differential forms on Lie algebroids over (pre-)multisymplectic manifolds
Noriaki Ikeda ***E-mail: nikeda@se.ritsumei.ac.jp
Department of Mathematical Sciences, Ritsumeikan University
Kusatsu, Shiga 525-8577, Japan
May 2, 2024
@@ 1 Introduction
A Poisson structure is an important geometric structure inspired by the Hamiltonian mechanics. A deformation of the Poisson structure by adding a closed $3$-form $H$ is called the twisted Poisson structure [30, 40]. We consider a bivector field $\pi\in\Gamma(\wedge^{2}TM)$ and a closed $3$-form $H\in\Omega^{3}(M)$ on a smooth manifold $M$. If $(\pi,H)$ satisfy
$\frac{1}{2}[\pi,\pi]_{S}=\langle\otimes^{3}\pi,\,H\rangle,$ (1)
$(\pi,H)$ is called a twisted Poisson structure, where $[-,-]_{S}$ is the Schouten bracket on the space of multivector fields and $\langle$-,\,$\rangle$ - , - ⟩ is the pairing of $TM$ and $T^{*}M$. The condition of the twisted Poisson structure is regarded as a compatibility condition between a Lie algebroid and a pre-$2$-plectic form. To consider a generalization to general pre-$n$-plectic form is natural. However a generalization is not straightforward since a simple generalization of $\pi$ and $H$ in Equation (1) to a multivector field and a pre-$n$-plectic form does not give a good condition.
A twisted Poisson structure was discovered in the analysis of a sigma model called a twisted Poisson (or a WZ Poisson) sigma model [30]. Recently, Chatzistavrakidis proposed a new higher dimensional topological sigma model with WZ term whose target space was a Poisson manifold [12, 13]. If and only if a multivector field on the target space is satisfied a geometric condition, the topological sigma model is consistent. A generalization to general Lie algebroid setting is formualated in the paper [26]. Consistency conditions of these sigma models imposes a geometric condition of an $E$-differential form with a Lie algebroid structure and a (pre-)(multi)symplectic form.. In this paper, we investigate this geometric structure mathematically.
Another motivation is a generalization of a momentum map on a symplectic manifold. A momentum map is defined on a symplectic manifold $M$ with a Lie group $G$ action. Let $\mathfrak{g}$ be a Lie algebra of the Lie group $G$. Recently, a generalization of a momentum map theory to a Lie algebroid setting has been proposed [5] inspired by analysis in [6, 33]. A momentum map is $\mu:M\rightarrow\mathfrak{g}^{*}$ is generalized to a section of the dual of a vector bundle $\mu\in\Gamma(E^{*})$ called a momentum section, where $E$ is a Lie algebroid. If a vector bundle is the direct product $E=M\times\mathfrak{g}$, it reduces to a momentum map.
@@
${}^{E}\mathrm{d}\mu=-\iota_{\rho}^{2}{\omega},$ (2)
@@
This paper is organized as follows. In Section 2, a Lie algebroid and related notion are prepared and notation is fixed. In Section 3, a compatible condition is defined and basic properties are analyzed. In Section 4, Examples are listed. In Section 5, the interpretation in terms of the higher Dirac structure is discussed. In Section 6, the compatible condition is formulated as a Q-manifold. In Section 7, we reformulate the compatible condition as a (twisted) QP-manifold. In the Appendix, some formulas are summarized.
@@ 2.1 Lie algebroids
Definition 2.1
Let $E$ be a vector bundle over a smooth manifold $M$. A Lie algebroid $(E,\rho,[-,-])$ is a vector bundle $E$ with a bundle map $\rho:E\rightarrow TM$ called the anchor map, and a Lie bracket $[-,-]:\Gamma(E)\times\Gamma(E)\rightarrow\Gamma(E)$ satisfying the Leibniz rule,
$[e_{1},fe_{2}]$ $=$ $f[e_{1},e_{2}]+\rho(e_{1})f\cdot e_{2},$ (3)
@@
Example 2.2 (Tangent Lie algebroids)
@@
Example 2.3 (Action Lie algebroids)
Suppose that there is a smooth action of a Lie group $G$ to a smooth manifold $M$, $M\times G\rightarrow M$. The differential of the map induces an infinitesimal action of the Lie algebra $\mathfrak{g}$ of $G$ on a manifold $M$. Since $\mathfrak{g}$ acts as a differential operator on $M$, the differential of the map determines a bundle map $\rho:M\times\mathfrak{g}\rightarrow TM$. Consistency of a Lie bracket requires that $\rho$ is a Lie algebra morphism such that
$~{}[\rho(e_{1}),\rho(e_{2})]$ $=$ $\rho([e_{1},e_{2}]),$ (4)
where the bracket in left hand side of (4) is a Lie bracket of vector fields. These data gives a Lie algebroid $(E=M\times\mathfrak{g},\rho,[-,-])$. This Lie algebroid is called an action Lie algebroid.
Example 2.4 (Poisson Lie algebroids)
A bivector field $\pi\in\Gamma(\wedge^{2}TM)$ is called a Poisson bivector field if $[\pi,\pi]_{S}=0$, where $[-,-]_{S}$ is a Schouten bracket on the space of multivector fields, $\Gamma(\wedge^{\bullet}TM)$. If a smooth manifold $M$ has a Poisson bivector field $\pi$, $(M,\pi)$ is called a Poisson manifold. For a Poisson bivector field $\pi$, a bundle map is defined as $\rho=-\pi^{\sharp}:T^{*}M\rightarrow TM$ by $\langle\pi^{\sharp}(\alpha),\,\beta\rangle=\pi(\alpha,\beta)$ for all $\beta\in\Omega^{1}(M)$.
Let $(M,\pi)$ be a Poisson manifold. Then, a Lie algebroid structure is introduced on $T^{*}M$. $\pi^{\sharp}:T^{*}M\rightarrow TM$ is the anchor map, and a Lie bracket on $\Omega^{1}(M)$ is defined by the so called Koszul bracket,
$[\alpha,\beta]_{\pi}={\mathcal{L}}_{\pi^{\sharp}(\alpha)}\beta-{\mathcal{L}}_{\pi^{\sharp}(\beta)}\alpha-\mathrm{d}(\pi(\alpha,\beta)),$ (5)
@@
Example 2.5 (Lie algebroids induced from twisted Poisson structures)
If a bivector field $\pi\in\Gamma(\wedge^{2}TM)$ and $H\in\Omega^{3}(M)$ satisfy
$\frac{1}{2}[\pi,\pi]_{S}=\langle\otimes^{3}\pi,\,H\rangle,$ (6)
$\mathrm{d}H=0,$ (7)
$(M,\pi,H)$ is called a twisted Poisson manifold [30, 36, 40, 27]. Here $\langle\otimes^{3}\pi,\,H\rangle$ is defined by $\langle\otimes^{3}\pi,\,H\rangle(\alpha_{1},\alpha_{2},\alpha_{3}):=H(\pi^{\sharp}(\alpha_{1}),\pi^{\sharp}(\alpha_{2}),\pi^{\sharp}(\alpha_{3}))$ for $\alpha_{i}\in\Omega^{1}(M)$.
If we take the same bundle map, $-\pi^{\sharp}:T^{*}M\rightarrow TM$ as in Example 2.4, and a Lie bracket deformed by $H$,
$[\alpha,\beta]_{\pi,H}={\mathcal{L}}_{\pi^{\sharp}(\alpha)}\beta-{\mathcal{L}}_{\pi^{\sharp}(\beta)}\alpha-\mathrm{d}(\pi(\alpha,\beta))+\iota_{\pi^{\sharp}(\alpha)}\iota_{\pi^{\sharp}(\beta)}H,$ (8)
for $\alpha,\beta\in\Omega^{1}(M)$. Then, $(T^{*}M,-\pi^{\sharp},[-,-]_{\pi,H})$ is a Lie algebroid.
@@ 2.2 Lie algebroid differentials
For a Lie algebroid $E$, sections of the exterior algebra of $E^{*}$ are called $E$-differential forms. Consider the spaces of $E$-differential forms $\Gamma(\wedge^{\bullet}E^{*})$. A differential ${}^{E}\mathrm{d}$ on $\Gamma(\wedge^{\bullet}E^{*})$ is called a Lie algebroid differential (or simply, a differential) ${}^{E}\mathrm{d}:\Gamma(\wedge^{m}E^{*})\rightarrow\Gamma(\wedge^{m+1}E^{*})$ and is defined as follows.
Definition 2.2
A Lie algebroid differential ${}^{E}\mathrm{d}:\Gamma(\wedge^{m}E^{*})\rightarrow\Gamma(\wedge^{m+1}E^{*})$ is defined by
${}^{E}\mathrm{d}\alpha(e_{1},\ldots,e_{m+1})$ $=$ $\sum_{i=1}^{m+1}(-1)^{i-1}\rho(e_{i})\alpha(e_{1},\ldots,\check{e_{i}},\ldots,e_{m+1})$ (9)
$+\sum_{1\leq i<j\leq m+1}(-1)^{i+j}\alpha([e_{i},e_{j}],e_{1},\ldots,\check{e_{i}},\ldots,\check{e_{j}},\ldots,e_{m+1}),$
@@
$({}^{E}\mathrm{d})^{2}=0$ is satisfied. The Lie algebroid differential ${}^{E}\mathrm{d}$ is a generalization of the de Rham differential on $T^{*}M$ and the Chevalley-Eilenberg differential on a Lie algebra.
3 Compatible $E$-$n$-forms
@@ 3.1 Definition and consistency
Definition 3.1
A closed $n+1$-form $\omega\in\Omega^{n+1}(M)$ is called a pre-$n$-plectic form. $(M,\omega)$ is called a pre-$n$-plectic manifold.
@@
Let ${{}^{E}\mathrm{d}}:\Gamma(\wedge^{m}E^{*})\rightarrow\Gamma(\wedge^{m+1}E^{*})$ be a Lie algebroid differential. Given $\omega\in\Omega^{n+1}(M)$, ${\iota_{\rho}^{k}}$ is defined by
$\iota_{\rho}^{k}{\omega}(v_{k+1},\ldots,v_{n+1})(e_{1},\ldots,e_{k})$ $=$ $\iota_{\rho(e_{1})}\ldots\iota_{\rho(e_{k})}{\omega}(v_{k+1},\ldots,v_{n+1})$ (10)
$=$ $\langle\otimes^{k}\rho,\,{\omega}\rangle(e_{k},\ldots,e_{1},v_{k+1},\ldots,v_{n+1})$
::=:= ${\omega}(\rho(e_{k}),\ldots,\rho(e_{1}),v_{k+1},\ldots,v_{n+1}),$
for $e_{1},\ldots,e_{k}\in\Gamma(E)$ and $v_{k+1},\ldots,v_{n+1}\in\mathfrak{X}(M)$.
Definition 3.2
@@
${}^{E}\mathrm{d}J=-\iota_{\rho}^{n+1}\omega=-\langle\otimes^{n+1}\rho,\,\omega\rangle.$ (11)
@@
${{}^{E}\mathrm{d}}(\iota^{n+1}_{\rho}\omega)=0,$ (12)
must be satisfied for consistency. The left hand side is
${{}^{E}\mathrm{d}}(\iota^{n+1}_{\rho}\omega)(e_{n+2},\ldots,e_{1})$
$=\mathrm{d}\omega(\rho(e_{1}),\ldots\rho(e_{n+2}))$
$+\sum_{1\leq i<j\leq n+2}(-1)^{i+j-1}\omega([\rho(e_{i}),\rho(e_{j})],\rho(e_{1}),\ldots,\check{\rho(e_{i})},\ldots,\check{\rho(e_{j})},\ldots,\rho(e_{n+2}))$
$+\sum_{1\leq i<j\leq n+2}(-1)^{i+j}\omega(\rho([e_{i},e_{j}]),\rho(e_{1}),\ldots,\check{\rho(e_{i})},\ldots,\check{\rho(e_{j})},\ldots,\rho(e_{n+2})),$ (13)
for $e_{i}\in\Gamma(E)$. Since ${}[\rho(e_{1}),\rho(e_{2})]=~{}\rho([e_{1},e_{2}])$ is satisfied in a Lie algebroid $E$, we obtain the following proposition.
Proposition 3.3
If $\omega$ is closed, Eq. (12) is satisfied and Equation (11) is consistent with $\rho$, [,][-,-][ - , - ] and $\omega$.

### v4

@@
[…]
Compatible $E$-Differential Forms on Lie Algebroids over (Pre-)Multisymplectic Manifoldss
[…]
Compatible $\bm{E}$-Differential Forms on Lie Algebroids
over (Pre-)Multisymplectic Manifolds
[…]
Noriaki IKEDA
[…]
N. Ikeda
[…]
Received November 13, 2023, in final form March 27, 2024; Published online March 31, 2024
[…]
@@
[…]
Poisson geometry; Lie algebroid; multisymplectic geometry; higher structures
[…]
53D17; 53D20; 58A50
@@ 1 Introduction
A Poisson structure is an important geometric structure inspired by the Hamiltonian mechanics. A deformation of the Poisson structure by adding a closed $3$-form $H$ is called the twisted Poisson structure [30, 40]. We consider a bivector field $\pi\in\Gamma\bigl{(}\wedge^{2}TM\bigr{)}$ and a closed $3$-form $H\in\Omega^{3}(M)$ on a smooth manifold $M$. If $(\pi,H)$ satisfy
$\frac{1}{2}[\pi,\pi]_{S}=\bigl{\langle}\otimes^{3}\pi,H\bigr{\rangle},$ (1.1)
$(\pi,H)$ is called a twisted Poisson structure, where $[-,-]_{S}$ is the Schouten bracket on the space of multivector fields and $\langle$-,\,$\rangle$ - , - ⟩ is the pairing of $TM$ and $T^{*}M$. The condition of the twisted Poisson structure is regarded as a compatibility condition between a Lie algebroid and a pre-$2$-plectic form. To consider a generalization to general pre-$n$-plectic form is natural. However, a generalization is not straightforward since a simple generalization of $\pi$ and $H$ in equation (1.1) to a multivector field and a pre-$n$-plectic form does not give a good condition.
A twisted Poisson structure was discovered in the analysis of a sigma model called a twisted Poisson (or a WZ Poisson) sigma model [30]. Recently, Chatzistavrakidis proposed a new higher dimensional topological sigma model with WZ term whose target space was a Poisson manifold [13, 14]. If and only if a multivector field on the target space is satisfied a geometric condition, the topological sigma model is consistent. A generalization to general Lie algebroid setting is formulated in the paper [25]. Consistency conditions of these sigma models imposes a geometric condition of an $E$-differential form with a Lie algebroid structure and a (pre-)(multi)symplectic form. In this paper, we investigate this geometric structure mathematically.
Another motivation is a generalization of a momentum map on a symplectic manifold. A momentum map is defined on a symplectic manifold $M$ with a Lie group $G$ action. Let $\mathfrak{g}$ be a Lie algebra of the Lie group $G$. Recently, a generalization of a momentum map theory to a Lie algebroid setting has been proposed [6] inspired by analysis in [5, 33]. A momentum map is ${\mu\colon M\rightarrow\mathfrak{g}^{*}}$ is generalized to a section of the dual of a vector bundle $\mu\in\Gamma(E^{*})$ called a momentum section, where $E$ is a Lie algebroid. If a vector bundle is the direct product $E=M\times\mathfrak{g}$, it reduces to a momentum map.
@@
${}^{E}\mathrm{d}\mu=-\iota_{\rho}^{2}{\omega},$
@@
This paper is organized as follows. In Section 2, a Lie algebroid and related notion are prepared and notation is fixed. In Section 3, a compatible condition is defined and basic properties are analyzed. In Section 4, examples are listed. In Section 5, the interpretation in terms of the higher Dirac structure is discussed. In Section 6, the compatible condition is formulated as a Q-manifold. In Section 7, we reformulate the compatible condition as a (twisted) QP-manifold.
@@ 2.1 Lie algebroids
Definition 2.1.
Let $E$ be a vector bundle over a smooth manifold $M$. A Lie algebroid $(E,\rho,[-,-])$ is a vector bundle $E$ with a bundle map $\rho\colon E\rightarrow TM$ called the anchor map, and a Lie bracket $[-,-]\colon\Gamma(E)\times\Gamma(E)\rightarrow\Gamma(E)$ satisfying the Leibniz rule,
$[e_{1},fe_{2}]=f[e_{1},e_{2}]+\rho(e_{1})f\cdot e_{2},$
@@
Example 2.3 (tangent Lie algebroids).
@@
Example 2.4 (action Lie algebroids).
Suppose that there is a smooth action of a Lie group $G$ to a smooth manifold $M$, $M\times G\rightarrow M$. The differential of the map induces an infinitesimal action of the Lie algebra $\mathfrak{g}$ of $G$ on a manifold $M$. Since $\mathfrak{g}$ acts as a differential operator on $M$, the differential of the map determines a bundle map $\rho\colon M\times\mathfrak{g}\rightarrow TM$. Consistency of a Lie bracket requires that $\rho$ is a Lie algebra morphism such that
$[\rho(e_{1}),\rho(e_{2})]=\rho([e_{1},e_{2}]),$ (2.1)
where the bracket in left-hand side of (2.1) is a Lie bracket of vector fields. These data gives a Lie algebroid $(E=M\times\mathfrak{g},\rho,[-,-])$. This Lie algebroid is called an action Lie algebroid.
Example 2.5 (Poisson Lie algebroids).
A bivector field $\pi\in\Gamma\bigl{(}\wedge^{2}TM\bigr{)}$ is called a Poisson bivector field if $[\pi,\pi]_{S}=0$, where $[-,-]_{S}$ is a Schouten bracket on the space of multivector fields, $\Gamma(\wedge^{\bullet}TM)$. If a smooth manifold $M$ has a Poisson bivector field $\pi$, $(M,\pi)$ is called a Poisson manifold. For a Poisson bivector field $\pi$, a bundle map is defined as $\rho=-\pi^{\sharp}\colon T^{*}M\rightarrow TM$ by ${\langle\pi^{\sharp}(\alpha),\,\beta\rangle=\pi(\alpha,\beta)}$ for all $\beta\in\Omega^{1}(M)$.
Let $(M,\pi)$ be a Poisson manifold. Then, a Lie algebroid structure is introduced on $T^{*}M$. $\pi^{\sharp}\colon T^{*}M\rightarrow TM$ is the anchor map, and a Lie bracket on $\Omega^{1}(M)$ is defined by the so called Koszul bracket,
$[\alpha,\beta]_{\pi}={\mathcal{L}}_{\pi^{\sharp}(\alpha)}\beta-{\mathcal{L}}_{\pi^{\sharp}(\beta)}\alpha-\mathrm{d}(\pi(\alpha,\beta)),$ (2.2)
@@
Example 2.6 (Lie algebroids induced from twisted Poisson structures).
If a bivector field $\pi\in\Gamma\bigl{(}\wedge^{2}TM\bigr{)}$ and $H\in\Omega^{3}(M)$ satisfy
$\frac{1}{2}[\pi,\pi]_{S}=\bigl{\langle}\otimes^{3}\pi,H\bigr{\rangle},\qquad\mathrm{d}H=0,$ (2.3)
$(M,\pi,H)$ is called a twisted Poisson manifold [27, 30, 36, 40]. Here $\bigl{\langle}\otimes^{3}\pi,H\bigr{\rangle}$ is defined by $\bigl{\langle}\otimes^{3}\pi,H\bigr{\rangle}(\alpha_{1},\alpha_{2},\alpha_{3}):=H\bigl{(}\pi^{\sharp}(\alpha_{1}),\pi^{\sharp}(\alpha_{2}),\pi^{\sharp}(\alpha_{3})\bigr{)}$ for $\alpha_{i}\in\Omega^{1}(M)$.
If we take the same bundle map, $-\pi^{\sharp}\colon T^{*}M\rightarrow TM$ as in Example 2.5, and a Lie bracket deformed by $H$,
$[\alpha,\beta]_{\pi,H}={\mathcal{L}}_{\pi^{\sharp}(\alpha)}\beta-{\mathcal{L}}_{\pi^{\sharp}(\beta)}\alpha-\mathrm{d}(\pi(\alpha,\beta))+\iota_{\pi^{\sharp}(\alpha)}\iota_{\pi^{\sharp}(\beta)}H$
for $\alpha,\beta\in\Omega^{1}(M)$. Then, $\bigl{(}T^{*}M,-\pi^{\sharp},[-,-]_{\pi,H}\bigr{)}$ is a Lie algebroid.
@@ 2.2 Lie algebroid differentials
For a Lie algebroid $E$, sections of the exterior algebra of $E^{*}$ are called $E$-differential forms. Consider the spaces of $E$-differential forms $\Gamma(\wedge^{\bullet}E^{*})$. A differential ${}^{E}\mathrm{d}$ on $\Gamma(\wedge^{\bullet}E^{*})$ is called a Lie algebroid differential (or simply, a differential) ${}^{E}\mathrm{d}\colon\Gamma(\wedge^{m}E^{*})\rightarrow\Gamma\bigl{(}\wedge^{m+1}E^{*}\bigr{)}$ and is defined as follows.
Definition 2.7.
A Lie algebroid differential ${}^{E}\mathrm{d}\colon\Gamma(\wedge^{m}E^{*})\rightarrow\Gamma\bigl{(}\wedge^{m+1}E^{*}\bigr{)}$ is defined by
${}^{E}\mathrm{d}\alpha(e_{1},\dots,e_{m+1})={}$ $\sum_{i=1}^{m+1}(-1)^{i-1}\rho(e_{i})\alpha(e_{1},\dots,\check{e_{i}},\dots,e_{m+1})$
$+\sum_{1\leq i<j\leq m+1}(-1)^{i+j}\alpha([e_{i},e_{j}],e_{1},\dots,\check{e_{i}},\dots,\check{e_{j}},\dots,e_{m+1}),$
@@
$\bigl{(}{}^{E}\mathrm{d}\bigr{)}^{2}=0$ is satisfied. The Lie algebroid differential ${}^{E}\mathrm{d}$ is a generalization of the de Rham differential on $T^{*}M$ and the Chevalley–Eilenberg differential on a Lie algebra.
3 Compatible $\bm{E}$-$\bm{n}$-forms
@@ 3.1 Definition and consistency
Definition 3.1.
A closed $(n+1)$-form $\omega\in\Omega^{n+1}(M)$ is called a pre-$n$-plectic form. $(M,\omega)$ is called a pre-$n$-plectic manifold.
@@
Let ${{}^{E}\mathrm{d}}\colon\Gamma(\wedge^{m}E^{*})\rightarrow\Gamma\bigl{(}\wedge^{m+1}E^{*}\bigr{)}$ be a Lie algebroid differential. Given $\omega\in\Omega^{n+1}(M)$, ${\iota_{\rho}^{k}}$ is defined by
$\iota_{\rho}^{k}{\omega}(v_{k+1},\dots,v_{n+1})(e_{1},\dots,e_{k})={}$ $\iota_{\rho(e_{1})}\dots\iota_{\rho(e_{k})}{\omega}(v_{k+1},\dots,v_{n+1})$
[…]
for $e_{1},\dots,e_{k}\in\Gamma(E)$ and $v_{k+1},\dots,v_{n+1}\in\mathfrak{X}(M)$.
Definition 3.2.
@@
${}^{E}\mathrm{d}J=-\iota_{\rho}^{n+1}\omega=-\bigl{\langle}\otimes^{n+1}\rho,\omega\bigr{\rangle}.$ (3.1)
@@
${{}^{E}\mathrm{d}}\bigl{(}\iota^{n+1}_{\rho}\omega\bigr{)}=0,$ (3.2)
must be satisfied for consistency. The left-hand side is
${{}^{E}\mathrm{d}}\bigl{(}\iota^{n+1}_{\rho}\omega\bigr{)}(e_{n+2},\dots,e_{1})=\mathrm{d}\omega(\rho(e_{1}),\dots,\rho(e_{n+2}))$
$\phantom{\qquad=}{}+\sum_{1\leq i<j\leq n+2}(-1)^{i+j-1}\omega([\rho(e_{i}),\rho(e_{j})],\rho(e_{1}),\dots,\check{\rho(e_{i})},\dots,\check{\rho(e_{j})},\dots,\rho(e_{n+2}))$
$\phantom{\qquad=}{}+\sum_{1\leq i<j\leq n+2}(-1)^{i+j}\omega(\rho([e_{i},e_{j}]),\rho(e_{1}),\dots,\check{\rho(e_{i})},\dots,\check{\rho(e_{j})},\dots,\rho(e_{n+2}))$
for $e_{i}\in\Gamma(E)$. Since $[\rho(e_{1}),\rho(e_{2})]=\rho([e_{1},e_{2}])$ is satisfied in a Lie algebroid $E$, we obtain the following proposition.
Proposition 3.3.
If $\omega$ is closed, (3.2) is satisfied and equation (3.1) is consistent with $\rho$, [,][-,-][ - , - ] and $\omega$.